OpenStudy (loser66):

is there any shortcut or general form to find roots of a complex number? like: find square root of i. Please, help

OpenStudy (loser66):

@dan815

OpenStudy (loser66):



OpenStudy (loser66):

the roots are \(\sqrt2/2+ i\sqrt 2/2 ~~and -\sqrt 2/2 -i\sqrt 2/2\) What is the shortcut of the procedure?

OpenStudy (mathmate):

Use De Moivre's theorem. The geometrical interpretation makes the calculations easy. try: http://www.mathamazement.com/Lessons/Pre-Calculus/06_Additional-Topics-in-Trigonometry/de-moivres-theorem.html

OpenStudy (loser66):

Thanks for the tip. I know how to find the roots out. Just wonder is there any shortcut but still be flawless to "cut" out the procedure. :)

OpenStudy (mathmate):

If you find an easier way, please share with me! xD

OpenStudy (loser66):

hehehe... sure.

OpenStudy (mathmate):

Thanks! :)